I'm totally confuse, please help!
Mary has a choice of using one ot two parking garages downtown:
Option 1: $8 an hour for the first two hours, then $2 an hour for each hour more than 2; or
Option 2: $4 for each hour
Let X= total hours parked. (I would like to use 6hrs, however I do not know how to set up the formula. I have to write a mathematical model representing the total parking cost, C, in terms of x for the following (assuming Mary is parking for more than 2 hours.
Option 1: C= 16
Option 2: C= 24

Answer by ad_alta(240)

I'll set up the models, and then you can decide which choice she should make based on how long she wants to park. If 't' is time in hours and 'C' is total cost, option 1 gives C=2*(t-2)+16 (t>=2) and option 2 gives C=4*t.